Masato is twice as old as Kim.Ten years ago Masato waas three times as old as Kim was ten years ago.How old is she now?
Answer by jorel1380(3719)

m=2k
m-10=3(k-10)
2k-10=3k-30
k=20
m=40
Masato is 40. Kim is 20..